Charles Lackey operates a bakery in Idaho​ Falls, Idaho. Because of its excellent product and excellent​ location, demand has increased by 35% in the last year. On far too many​ occasions, customers have not been able to purchase the bread of their choice. Because of the size of the​ store, no new ovens can be added. At a staff​ meeting, one employee suggested ways to load the ovens differently so that more loaves of bread can be baked at one time. This new process will require that the ovens be loaded by​ hand, requiring additional manpower. This is the only thing to be changed.  The bakery currently makes 1,500 loaves per month. The pay will be ​$88 per hour for employees and each employee works 160 hours per month.

Charles Lackey can also improve the yield by purchasing a new blender. The new blender will mean an increase in his investment. This new blender will mean an increase in his costs of ​$125 per​ month, but he will achieve the same new output​ (an increase to 2,025 as the change in labor hours.

A) Current productivity for 640 work hours​ =.293 ​loaves/dollar

If Charles chooses to increase the number of work hours to 864 in order to employ the new oven loading​ technique, then the productivity is​ =.293 ​loaves/dollar ​

B) If Charles instead chooses to purchase a new blender​ (while holding labor constant at 640 hours at $8 per ​hour)​,then the productivity is​ =.386 ​loaves/dollar

C) By adding​ manpower, the percentage increase in productivity is __________
Answer #1

A) Current productivity for 640 work hours​ =1500/(640*8) = .293​loaves/dollar

If Charles chooses to increase the number of work hours to 864 in order to employ the new oven loading​ technique, then the productivity is​ =(1500*1.35)/(864*8) = .293 ​loaves/dollar ​

B) If Charles instead chooses to purchase a new blender​ (while holding labor constant at 640 hours at $8 per ​hour)​,then the productivity is​ =(1500*1.35)/(640*8+125)=0.386 ​loaves/dollar

C) By adding​ manpower, the percentage increase in productivity is 0% [productivity remained same]